OpenStudy (anonymous):

Given ~ and AB = 20, CA = 8, and US = 9, what is ST?

OpenStudy (anonymous):

ABC~STU

OpenStudy (anonymous):

@amistre64 @welshfella

OpenStudy (igreen):

\(\sf \color{red}{AB} C = \color{red}{ST}U\)

OpenStudy (igreen):

So what's our answer going to be?

OpenStudy (anonymous):

20 but people kept saying it was 22.5

OpenStudy (igreen):

20 is correct.

OpenStudy (anonymous):

http://brainly.com/question/284332

OpenStudy (igreen):

It shows this, right? ≅

OpenStudy (igreen):

ABC ≅ STU

OpenStudy (anonymous):

yes the equal to sign

OpenStudy (igreen):

Oh..I think he's right.

OpenStudy (timbo5lice):

lol

OpenStudy (igreen):

Yes..we set up a proportion.. US and CA are similar to each other. So there's a proportion. \(\sf \dfrac{9}{8} = \dfrac{x}{20}\) Cross multiply

OpenStudy (anonymous):

180/8

OpenStudy (igreen):

8 * x = 8x 9 * 20 = 180 \(\sf 8x = 180\)

OpenStudy (igreen):

Yep. \(\sf x = \dfrac{180}{8}\)